OpenStudy (anonymous):

x+2/x+3 Name the vertical asymptotes, if any Name the horizontal asymptotes, if any

OpenStudy (anonymous):

Vertical asymptotes are given by the domain. In your case, it would be -3 because you can't divide by 0. And horizontal asymptotes (also oblicuous) in this functions are given by the first coefficient, both top and down. In this case, it would be 1/1, which is 1, which means you have an horizontal asymptote on Y=1 , and a vertical on X=-3

OpenStudy (anonymous):

What? No that's not true. VA: x=-3 HA: y=0 because your HA is finding x for your denominator and y is your y-int. you do not have aa y-int. in this problem so y=0.

OpenStudy (anonymous):

?? Dude, 0 isn't a HA. 1 is a HA because if it wasn't, then x+3 = x+2 , which is absurd...

OpenStudy (anonymous):

NO. If it was x+2/x+3 + 1 then you would be right, but it's not. x+2/x is your parent function. x+3 is how you get VA you get HA by seeing whatever is added or subtracted to your x-int. which would be your y-int. y=0 because there's nothing added or subtracted in your y-int.

OpenStudy (anonymous):

Please man, go read about homographic functions first. If not, go learn about something called limits, and see what happens with the function in infinite. There u have your definition of Horizontal Asymptote, and then come back and tell me that the horizontal asymptote is still 0..... i even explained you that x+3 != x+2, and therefore Y can never be 1 .....

OpenStudy (anonymous):

Okay I'm in Honors Algebra 2; I just took a quiz on this Monday and made a 100 so don't you think I would know what I'm doing.

OpenStudy (anonymous):

Haha, i'm just telling you that you are wrong. I even explained how Horizontal Asymptotes work. What does it matter if you are in honors Algebra 2 (wtf is that anyway?) or in highscool?

OpenStudy (anonymous):

(x+2)/(x+3) VA= when bottom = 0 so x+3=0 -> is -3 H= where your x is very very large. just pick a very large x and plug it into x. which is 1

OpenStudy (anonymous):